OpenStudy (albert0898):

The Fg between 2 masses is 25N. If both masses are doubled and the distance is doubled (all at the same time), what is the new Fg? Use the formula: F= Gm1m2/r

OpenStudy (aum):

Are you sure the formula for F is correct? Shouldn't the denominator be r^2?

OpenStudy (albert0898):

It should be r^2

OpenStudy (aum):

\[ \large F_g = \frac{Gm_1m_2}{r^2} \\ 25 = \frac{Gm_1m_2}{r^2} ~~ \text{ ----- (1)} \\ \text{Both masses are doubled and the distance is doubled too:} \\ F_g = \frac{G(2m_1)(2m_2)}{(2r)^2} = \frac{4Gm_1m_2}{4r^2} = \frac{Gm_1m_2}{r^2} = 25~\text{N from (1)} \\ \]

OpenStudy (aum):

So the force remains the same at 25 N.

OpenStudy (albert0898):

Thank you :)